OpenStudy (anonymous):

During a long drive, Tony counted the number of pickups and sedans he saw driving in the opposite direction. After a while, he noticed that on average there were 5 sedans for every 2 pickups. At this ratio, how many sedans would he have counted if he had passed 18 pickups?

OpenStudy (loser66):

to this kind of problem, you should make up a ratio, like this. \(\dfrac{sedan}{pickup}=\dfrac{5}{2}\)remember, this is sedan in above. or \(\dfrac{pickup}{sedan}=\dfrac{2}{5}\) Pay attention, pickup is in above. Look at your question, what is they ask you? if they ask for sedan, so, the ratio you should pick to solve will have pickup at above. that's is 2/5 , then time it with number of number of pickup they give you. and vice versa. I mean